OpenStudy (eriyalovesu):

Choose the equation below that represents the line that passes through the point (2, 4) and has a slope of 3. y − 4 = 3(x − 2) y − 2 = 3(x − 4) y + 4 = 3(x + 2) y + 2 = 3(x + 4)

pooja195 (pooja195):

Seems that they want the answer in point slope form: \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (eriyalovesu):

y=mx+b

OpenStudy (eriyalovesu):

oh point slope from

pooja195 (pooja195):

Thats slope intercept your options are asking for point slope form

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 Seems that they want the answer in point slope form: \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1) \(\color{#0cbb34}{\text{End of Quote}}\) Use this info to plug in your given values

OpenStudy (eriyalovesu):

so like 2-4=3(x-2)

pooja195 (pooja195):

y remains the same y1 changes so that would be incorrect Givens: m=3 x1=2 y1=4 Plug those in :)

OpenStudy (eriyalovesu):

oh! Its A y-4=3(x-2)

pooja195 (pooja195):

Perfect! :D

OpenStudy (eriyalovesu):

Thank you!